# Parameters vs FLOPs: Scaling Laws for Optimal Sparsity for Mixture-of-Experts Language Models

### v2

@@ 8 Conclusion
MoEs were originally introduced to allow increasing model capacity without a significant increase in inference cost. Our experiments show that under fixed total training compute budget increasing sparsity in MoEs leads to smaller FLOPs per example, higher number of parameters, and lower pretraining loss simultaneously. In other words, in the context of MoEs, if there are no constraints on the total number of parameters, increasing the capacity of the model through parameter count seem to be the optimal strategy if lower pretraining loss is the main goal. On the other hand, when comparing how well the pretraining performance transfers to various downstream tasks, denser models exhibit better transfer performance on certain types of task that potentially rely on deeper processing of the input vs the knowledge stored in the parameters of the model. This potentially signals the importance of the role of FLOPs per example in increasing the capacity of the model during inference. This observation reveals an interesting direction to improve the performance efficiency of MoEs at inference time.
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Experimental Setup
We train and evaluate auto-regressive sparse Mixture-of-Experts (MoE) language models of varying sizes and configurations on subsets of the RedPajamaV1 dataset (together2023redpajama). The key variables we explore in our experiments are total model parameters $N$, training compute budget $C$, and the MoE sparsity $S$.
@@
Our models are pre-trained on subsets of the RedPajamaV1 dataset [^4] GitHub repository: https://github.com/togethercomputer/RedPajama-Data (together2023redpajama), which attempts to replicate the LLaMA pre-training data recipe and comprises 1.2 trillion tokens from sources such as Common Crawl, C4, GitHub, and Wikipedia. In all our experiments, the effective dataset size is adjusted based on the training compute budget $C$ and the model size $N$. We tokenize the data using the GPT-NeoX tokenizer (black2022gpt), which has a vocabulary size of $50,432$ tokens.
@@
We use auto-regressive transformer-based MoE language models in order to study compute-parameter trade-offs by varying MoE sparsity. We use the Megablocks library (megablocks) to train dropless MoEs in which the routing mechanism ensures that all tokens are efficiently routed without being dropped due to routing capacity constraints.
@@
We fix a subset of hyperparameters for which changing values in preliminary experiments (a) did not significantly improve pre-training loss, (b) the optimal value remained the same across several model configurations, or (c) in order to reduce the search space (i.e., limited compute resources). Specifically, we first opted to use $z$-router loss (zoph2022st) and $qk$-normalization (wortsman2023small) in order to stabilize training for large MoEs. Second, we fixed MoE router jitter noise to 00, as it did not improve performance. We also fixed our batch size to $1024$ for all model sizes.
We swept over hyperparameters that, when adjusted, (a) significantly improved pre-training loss and (b) the optimal values varied across different model configurations. We increase the MoE sparsity by decreasing the number of active experts and/or increasing the number of total experts. We also varied the MoE granularity (ludziejewski2024scaling), MoE load balancing regularizer, Adam learning rate, and linear warm-up steps (fraction) in order to improve pre-training loss. The table below summarizes our hyperparameter sweeps:
@@ Appendix C Estimating Mixture-of-Expert (MoE) FLOPs
Similar to prior work on scaling laws (e.g.,  kaplan_scale; hoffmann2022training; ludziejewski2024scaling), we use theoretical FLOP estimates as proxies for training and inference costs of language models. In this section, we (a) outline our methodology for estimating FLOPs for MoEs and (b) show that the proposed estimator closely approximates empirical FLOPs of large-scale MoEs.
@@
Consider an MoE model with $n_{\text{layers}}$ MoE layers, each with an embedding dimension of $d_{\text{model}}$. We denote the number of total experts and active experts in each MoE layer by $E_{\text{total}}$ and $E_{\text{active}}$ respectively. Following ludziejewski2024scaling, we let $G$ denote the MoE granularity, which defaults to $1$ and controls the size of each expert relative to the size of a feed-forward layer in an equivalent dense transformer. In order to change sparsity in a more granular manner, we treat the number of active experts as an independent variable that does not scale with granularity $G$. In our experiments, we use a vocabulary size $n_{\text{vocab}}=50,432$, a context length $n_{\text{ctx}}$ of $2048$, and GLU modules (Gated Linear Units) (shazeer2017) over feed-forward modules as the architecture of choice for MoE experts. We also set the (a) hidden dimension of each GLU expert $d_{\text{ffn}}$ to $4\cdot d_{\text{model}}$ and (b) instantiate MoEs where the number of attention heads $n_{\text{heads}}$ times the dimensionality for each head $d_{\text{head}}$ equals $d_{\text{model}}$, i.e., $n_{\text{heads}}d_{\text{head}}=d_{\text{model}}$.
@@
Like in prior work on scaling laws (kaplan_scale; hoffmann2022training), we take a two-step approach to estimate module-specific FLOPs. Given a module, we first estimate the number of parameters in the module and then scale this with an appropriate constant corresponding to the number of add-multiply operations per parameter through a forward and backward pass of the given module. We also omit non-leading terms such as non-linearities, biases, and layer normalization in our estimation. We estimate the FLOPs per token for attention modules, MoE routers, MoE experts, and the final un-embedding layer as follows:
@@
    • QKV (and final) projections. These projections involve $4\cdot d_{\text{model}}n_{\text{heads}}d_{\text{heads}}=4d^{2}_{\text{model}}$ parameters. Following kaplan_scale, we use the multiplicative constant $C=6$ to account for the add-multiply operations per parameter in a forward and backward pass through linear modules, resulting in a FLOPs-per-token estimate of $4\cdot C\cdot d^{2}_{\text{model}}$.
@@
    • Router. The MoE routing linearly maps a $d_{\text{model}}$-dimensional token embedding to a $E_{\text{total}}$-dimensional logit vector, which is subsequently used to map the token to $E_{\text{active}}$ active experts. Following ludziejewski2024scaling, we use a multiplicative constant $R=14$ that accounts for the add-multiply-route operations per router parameter. The resulting FLOP estimate equals $R\cdot d_{\text{model}}E_{\text{total}}$
    • Experts. Each MoE experts corresponds to a GLU module (shazeer2017) with $d_{\text{ffn}}=4\cdot d_{\text{model}}$. Since there are $E_{\text{active}}$ active experts with granularity $G$, each involving three linear projections, this results in a FLOP estimate of $\nicefrac{{1}}{{G}}\cdot 3\cdot E_{\text{active}}\cdot C\cdot d_{\text{model}}d_{\text{ffn}}=\nicefrac{{12C}}{{G}}\cdot E_{\text{active}}\cdot d^{2}_{\text{model}}$.
@@
For standard dense transformers, the FLOPs are often estimated as $6ND$ (kaplan_scale; hoffmann2022training). Given that $D$ is fixed and not adjusted dynamically, $N$ can serve as a relative estimator of FLOPs per token for dense transformer models.
@@ D.4 Comparing IsoFLOP Surface Analysis with Independent 2d IsoFLOPs
In this section, we evaluate whether these findings remain consistent when we do not rely on fitted IsoFLOP surfaces. Specifically, similar to Approach II in hoffmann2022training, we directly fit univariate quadratic functions that map model size $N$ to pre-training loss $L$, independently for each sparsity level and training compute budget. We then assess these univariate fits to determine whether our findings in Section 2 hold.
@@
Appendix E Incorporating Sparsity into Scaling Laws
@@
Table 2 shows the parameters used to initialize L-BFGS used to fit the proposed parametric scaling law given in Equation 6. Table 3 shows the estimated parameters for the parameteric model. We use a held out dataset that consists of data points for models with sparsity value $S=0.98$ to validate the performance of the estimated model coefficients. The mean squared error and the Huber loss error on the dataset used to fit the model is $0.00056$ and $0.0036$ respectively and $0.0058$ and $0.0011$ respectively on the out-of-sample validation set.

### v3

@@ 8 Conclusion
MoEs were originally introduced to allow increasing model capacity without a significant increase in inference cost. Our experiments show that under fixed total training compute budget increasing sparsity in MoEs leads to smaller FLOPs per example, higher number of parameters, and lower pretraining loss simultaneously. In other words, in the context of MoEs, if there are no constraints on the total number of parameters, increasing the capacity of the model through parameter count seem to be the optimal strategy if lower pretraining loss is the main goal. On the other hand, when comparing how well the pretraining performance transfers to various downstream tasks, denser models exhibit better transfer performance on certain types of task that potentially rely on deeper processing of the input vs the knowledge stored in the parameters of the model. This potentially signals the importance of the role of FLOPs per example in increasing the capacity of the model during inference. Our experiments demonstrate that MoEs use Chain-of-Thought prompting more effectively than dense models, achieving better performance when allocated additional computational resources during inference. This observation reveals an interesting direction to improve the performance efficiency of MoEs at inference time.
@@
Acknowledgments
The authors would like to thank Vaishaal Shankar, Fartash Faghri, Skyler Seto, Mustafa Shukor, Amitis Shidani, David Grangier, Etai Littwin, Alexander Toshev and Preetum Nakkiran for their insightful discussions, feedback and technical support that significantly contributed to the development of this paper.
@@ Appendix B Experimental Setup
We train and evaluate auto-regressive sparse Mixture-of-Experts (MoE) language models of varying sizes and configurations on subsets of the RedPajamaV1 dataset Together Computer (2023). The key variables we explore in our experiments are total model parameters $N$, training compute budget $C$, and the MoE sparsity $S$.
@@
Our models are pre-trained on subsets of the RedPajamaV1 dataset [^4] GitHub repository: https://github.com/togethercomputer/RedPajama-Data Together Computer (2023), which attempts to replicate the LLaMA pre-training data recipe and comprises 1.2 trillion tokens from sources such as Common Crawl, C4, GitHub, and Wikipedia. In all our experiments, the effective dataset size is adjusted based on the training compute budget $C$ and the model size $N$. We tokenize the data using the GPT-NeoX tokenizer Black et al. (2022), which has a vocabulary size of $50,432$ tokens.
@@
We use auto-regressive transformer-based MoE language models in order to study compute-parameter trade-offs by varying MoE sparsity. We use the Megablocks library Gale et al. (2023) to train dropless MoEs in which the routing mechanism ensures that all tokens are efficiently routed without being dropped due to routing capacity constraints.
@@
We followed established best practices to train MoEs that included carefully searching over important hyperparameters like learning rate, weight decay, warm up schedule. Furthermore, we used a load balancing loss, router-Z loss to stabilize training and QK-normalization to stabilize training. We fix a subset of hyperparameters for which changing values in preliminary experiments (a) did not significantly improve pre-training loss, (b) the optimal value remained the same across several model configurations, or (c) in order to reduce the search space (i.e., limited compute resources). Specifically, we first opted to use $z$-router loss Zoph et al. (2022) and $qk$-normalization Wortsman et al. (2023) in order to stabilize training for large MoEs. Second, we fixed MoE router jitter noise to 00, as it did not improve performance. We also fixed our batch size to $1024$ for all model sizes.
We swept over hyperparameters that, when adjusted, (a) significantly improved pre-training loss and (b) the optimal values varied across different model configurations. We increase the MoE sparsity by decreasing the number of active experts and/or increasing the number of total experts. We also varied the MoE granularity Ludziejewski et al. (2024), MoE load balancing regularizer, Adam learning rate, and linear warm-up steps (fraction) in order to improve pre-training loss. The table below summarizes our hyperparameter sweeps:
@@ Appendix C Estimating Mixture-of-Expert (MoE) FLOPs
Similar to prior work on scaling laws (e.g.,  Kaplan et al. (2020); Hoffmann et al. (2022); Ludziejewski et al. (2024)), we use theoretical FLOP estimates as proxies for training and inference costs of language models. In this section, we (a) outline our methodology for estimating FLOPs for MoEs and (b) show that the proposed estimator closely approximates empirical FLOPs of large-scale MoEs.
@@
Consider an MoE model with $n_{\text{layers}}$ MoE layers, each with an embedding dimension of $d_{\text{model}}$. We denote the number of total experts and active experts in each MoE layer by $E_{\text{total}}$ and $E_{\text{active}}$ respectively. Following Ludziejewski et al. (2024), we let $G$ denote the MoE granularity, which defaults to $1$ and controls the size of each expert relative to the size of a feed-forward layer in an equivalent dense transformer. In order to change sparsity in a more granular manner, we treat the number of active experts as an independent variable that does not scale with granularity $G$. In our experiments, we use a vocabulary size $n_{\text{vocab}}=50,432$, a context length $n_{\text{ctx}}$ of $2048$, and GLU modules (Gated Linear Units) (Shazeer et al., 2017) over feed-forward modules as the architecture of choice for MoE experts. We also set the (a) hidden dimension of each GLU expert $d_{\text{ffn}}$ to $4\cdot d_{\text{model}}$ and (b) instantiate MoEs where the number of attention heads $n_{\text{heads}}$ times the dimensionality for each head $d_{\text{head}}$ equals $d_{\text{model}}$, i.e., $n_{\text{heads}}d_{\text{head}}=d_{\text{model}}$.
@@
Like in prior work on scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), we take a two-step approach to estimate module-specific FLOPs. Given a module, we first estimate the number of parameters in the module and then scale this with an appropriate constant corresponding to the number of add-multiply operations per parameter through a forward and backward pass of the given module. We also omit non-leading terms such as non-linearities, biases, and layer normalization in our estimation. We estimate the FLOPs per token for attention modules, MoE routers, MoE experts, and the final un-embedding layer as follows:
@@
    • QKV (and final) projections. These projections involve $4\cdot d_{\text{model}}n_{\text{heads}}d_{\text{heads}}=4d^{2}_{\text{model}}$ parameters. Following Kaplan et al. (2020), we use the multiplicative constant $C=6$ to account for the add-multiply operations per parameter in a forward and backward pass through linear modules, resulting in a FLOPs-per-token estimate of $4\cdot C\cdot d^{2}_{\text{model}}$.
@@
    • Router. The MoE routing linearly maps a $d_{\text{model}}$-dimensional token embedding to a $E_{\text{total}}$-dimensional logit vector, which is subsequently used to map the token to $E_{\text{active}}$ active experts. Following Ludziejewski et al. (2024), we use a multiplicative constant $R=14$ that accounts for the add-multiply-route operations per router parameter. The resulting FLOP estimate equals $R\cdot d_{\text{model}}E_{\text{total}}$
    • Experts. Each MoE experts corresponds to a GLU module (Shazeer et al., 2017) with $d_{\text{ffn}}=4\cdot d_{\text{model}}$. Since there are $E_{\text{active}}$ active experts with granularity $G$, each involving three linear projections, this results in a FLOP estimate of $\nicefrac{{1}}{{G}}\cdot 3\cdot E_{\text{active}}\cdot C\cdot d_{\text{model}}d_{\text{ffn}}=\nicefrac{{12C}}{{G}}\cdot E_{\text{active}}\cdot d^{2}_{\text{model}}$.
@@
For standard dense transformers, the FLOPs are often estimated as $6ND$ (Kaplan et al., 2020; Hoffmann et al., 2022). Given that $D$ is fixed and not adjusted dynamically, $N$ can serve as a relative estimator of FLOPs per token for dense transformer models.
@@ D.4 Comparing IsoFLOP Surface Analysis with Independent 2d IsoFLOPs
In this section, we evaluate whether these findings remain consistent when we do not rely on fitted IsoFLOP surfaces. Specifically, similar to Approach II in Hoffmann et al. (2022), we directly fit univariate quadratic functions that map model size $N$ to pre-training loss $L$, independently for each sparsity level and training compute budget. We then assess these univariate fits to determine whether our findings in Section 2 hold.
@@
Appendix E Does Chain-of-Thought prompting benefit sparse MoEs more than dense models?
In Section 4, we observed that dense models fare marginally better than sparse MoEs on reading comprehension tasks, potentially due to the higher inference-time compute of a dense model than a perplexity-matched sparse MoE. Then, in Section 6, we hypothesized that alternative strategies to increase inference-time compute may reduce the gap between sparse MoEs and dense models on such tasks. In this section, we test this hypothesis by leveraging a “length-controlled” variant of few-shot Chain-of-Thought (CoT) prompting to indirectly control inference-time compute. We then use this to study the effect of inference-time compute on downstream task performance of dense and MoE models.
Experiment setup.
We evaluate Qwen1.5 models (Bai et al., 2023) on the GSM8k dataset (Cobbe et al., 2021) to study the effect of few-shot CoT prompting (Wei et al., 2022b) on downstream task performance. We look at the effect of increasing inference-time compute (via CoT prompting) on GSM8k performance of dense models with sizes ranging from 0.5B to 14B and a 5x2.7B sparse MoE. We also use $10$ fixed examples from the GSM8k train split as few-shot examples for all runs.
Length-controlled CoT prompting enables control over inference-time compute.
Given an instruction-tuned model and a problem from the GSM8k dataset, we control the inference-time compute of the model by controlling the number of tokens generated to output the final answer to the given problem. To do so, we observe that providing instructions via system prompts (with few-shot CoT prompting) does not effectively control the number of generated tokens and, as a result, inference-time compute. We also observe that the average number of tokens in the few-shot GSM8k answers (provided in-context) strongly influences the number of tokens generated by the model to solve the given GSM8k question. Therefore, similar to work on designing GSM8k variants to analyze language modeling phenomena (Mirzadeh et al., 2024; Li et al., 2024; Zhang et al., 2024), we prompt an instruction-tuned model—Llama-3.1-70b (Dubey et al., 2024) in our experiment—to rewrite GSM8k answers in approximately $k\in\{5,\dots,100\}$ words. Then, we use the paraphrased GSM8k examples as few-shot examples to indirectly control the number of generated tokens. As shown in Figure 13(a), this approach enables systematic control over the length of paraphrased answers. The strong correlation ($\rho=0.88$) between few-shot and generated answer lengths in Figure 13(b) validates that our approach effectively modulates inference-time compute.
Figure 13: Varying inference-time compute via length-controlled Chain-of-Thought prompting. We can control inference-time compute (number of generated tokens) via Chain-of-Thought prompting in two steps: (a) generating paraphrased GSM8k answers of varying lengths (5-100 words) and (b) using paraphrased answers as few-shot examples in CoT prompts to influence the length of generated answers determine the model’s output length ($\rho=0.88$ correlation). The systematic shift in answer length distributions in subplot (a) and the linear relationship between few-shot answer length and generated answer lenth ($\rho=0.88$ correlation) in (b) validate that our prompting approach effectively modulates inference-time compute, enabling controlled studies of its impact on model performance.
Effect of length-controlled CoT on downstream task performance.
We investigate how inference-time compute affects GSM8k performance across Qwen1.5 models using $10$-shot length-controlled CoT prompting, varying target answer lengths $k$ from $1$ to roughly $70$ words on average. As shown in Figure 14, when we increase inference-time compute indirectly through longer generated answers ($x$-axis), we observe a roughly linear improvement in GSM8k performance ($y$-axis) across all model sizes. The linear fits for each model also indicate a pattern through their slopes $m$: larger dense models benefit more from additional inference-time compute. This effect is particularly striking when comparing models at the extremes (i.e., Qwen1.5-0.5B versus Qwen1.5-14B).
To analyze whether inference-time compute affects dense and sparse MoE models differently, we examine how the “performance” slope $m$ (i.e., accuracy gain per generated word) varies with model size. While the Qwen1.5-5x2.7B MoE cannot be directly compared to dense models due to different active parameter counts, we can account for this by plotting $m$ against the number of active parameters for both architectures. As shown in Figure 15, when controlling for active parameter count, the MoE model exhibits a higher slope than would be expected from interpolating between dense models. This suggests that MoE models benefit more from dynamically increased inference-time compute compared to dense models with equivalent active parameters.
Figure 14: Effect of length-controlled CoT prompting on GSM8k performance across model scales. We evaluate the relationship between inference-time compute (controlled via answer length) and GSM8k accuracy for dense Qwen1.5 models (0.5B-14B parameters) and a 5x2.7B sparse MoE. For all models, increased inference-time compute improves accuracy roughly linearly, with slopes $m$ indicating the marginal effect. Larger dense models show steeper slopes, demonstrating they benefit more from additional inference-time compute—for example, the 14B model’s performance improves from 20% to 70% as answer length increases, while the 0.5B model improves only from 1% to 5%.
Figure 15: Sparse MoEs benefit more from increased inference-time compute than dense models. We plot the marginal effect of inference-time compute (accuracy gain per generated word) against model size for dense Qwen1.5 models and a 5x2.7B sparse Qwen1.5 MoE. When controlling for the number of active parameters, the MoE model (orange) shows a higher performance slope than would be expected from interpolating between dense models (blue), suggesting that sparse MoEs benefit more from dynamically increased inference-time compute via strategies like CoT prompting.
Appendix F Incorporating Sparsity into Scaling Laws
@@
Table 2 shows the parameters used to initialize L-BFGS used to fit the proposed parametric scaling law given in Equation 6. Table 3 shows the estimated parameters for the parameteric model. We use a held out dataset that consists of data points for models with sparsity value $S=0.98$ to validate the performance of the estimated model coefficients. The mean squared error and the Huber loss error on the dataset used to fit the model is $0.00056$ and $0.0036$ respectively and $0.0058$ and $0.0011$ respectively on the out-of-sample validation set. The quality of the fit measured via the $R^{2}$ metric is $99\%$ on fitting data and $68\%$ on the held out validation dataset.